OpenStudy (anonymous):

what is 24/35 divided by 6?

OpenStudy (anonymous):

24/35 x 35/210

OpenStudy (anonymous):

please show steps that way i can complete the work thanks

OpenStudy (anonymous):

24/35 divided by 6 = 24/35x 1/6 =24/210

OpenStudy (anonymous):

24/35 / 6 = 24/35 * 1/6 (you just flip the 6). multiply the denoms and numes. (24*1)/(35*6)= 24/210

OpenStudy (anonymous):

.1175

OpenStudy (anonymous):

840/7350

OpenStudy (anonymous):

I agree with donny55 dont know about the other answers